The segment [a,b],where A(-3,1) and B(5,7) is a diameter of a circle c,
Find the coordinates of the centre and calculate the length of the radius.

Answer by stanbon(75887)

The segment [a,b],where A(-3,1) and B(5,7) is a diameter of a circle c,
Find the coordinates of the centre and calculate the length of the radius.
------
center = [(-3+5)/2 , (1+7)/2] = (1,4)
------------------
radius = sqrt[(5-1)^2 + (7-4)^2] = sqrt[16+9] = 5
---------------------------
Cheers,
Stan H.
-------------------